# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ A.1 Proof of Lemma 1
We prove the equivalence statement by contradiction. Suppose the Vandermonde matrix $V_{A}$ is not invertible. Because $V_{A}$ is a square matrix, there exists a vector $b\neq 0$, such that $V_{A}b=0$. Assume furthermore that there exists a polynomial $P(x)$ satisfying ${P}(w_{i})=y_{i},\forall 1\leq i\leq m$, otherwise the statement holds. Then the polynomial $\tilde{P}(x)=P(x)+\sum_{i=1}^{m}b[i]x^{k_{(i)}}$ also satisfies $\tilde{P}(w_{i})=y_{i},\forall 1\leq i\leq m$, hence in view of $b\neq 0$, there exist at least two solutions to the interpolation problem.
Next assume that the Vandermonde matrix $V_{A}$ is invertible. Then equation (16) provides a solution. We prove by contradiction that this is the unique solution. Assume that there are at least two polynomials $P_{1}\neq P_{2}$ satisfying $P_{j}(w_{i})=y_{i},j=1,2$, then $P_{1}-P_{2}\neq 0$ and $(P_{1}-P_{2})(w_{i})=0,\forall 1\leq i\leq m$. Writing the difference in the form $(P_{1}-P_{2})(x)=\sum_{i=1}^{m}b[i]x^{k_{(i)}}$ and introducing the notation $b=(b[1],b[2],\cdots,b[m])^{T}\in\mathbb{R}^{m}$, $b\neq 0$, we get $V_{A}b=0$, which contradicts with the assumption that $V_{A}$ is invertible.
A.2 Proof of Lemma 2
@@
We first prove (2) $\Rightarrow$ (1). Since the smallest singular value of the Vandermonde matrix $V_{A}$ is positive, $V_{A}$ is invertible. In view of Lemma 1, the set $A$ is unisolvent with respect to $\mathscr{P}_{l}(\mathbb{R}^{d})$ and hence also unisolvent for $\mathscr{P}_{l}(\Omega)$. Therefore, all polynomials $P\in\mathscr{P}_{l}(\mathbb{R}^{d})$ can be written in the form
$P(x)=P(A)^{T}V_{A}^{-1}v_{x}.$ (41)
@@
We proceed to prove (1) $\Rightarrow$ (2). First we prove by contradiction that the Vandermonde matrix $V_{A}$ is invertible. Assume that this is not true. Then there exists a non-zero vector $b\in\mathbb{R}^{m}$, such that $V_{A}^{T}b=0$. Then the corresponding polynomial $P(x)=b^{T}v_{x}\in\mathcal{P}_{l}(\Omega)$ satisfies $P(x)=0$, $\forall x\in A$, i.e. $\|P\|_{A}=0$. However, since $b\neq 0$, $\exists x^{\prime}\in\Omega$ such that $P(x^{\prime})\neq 0$, i.e. $\sup_{x}|P(x)|>0$. This contradicts to the norming set assumption of $A$.
To establish $c_{S}^{-1}<C_{0}c_{N}$, we need to introduce a “standard” norming set $A^{*}$ as a “bridge” between $c_{S}$ and $c_{N}$. Define the set $A^{*}$ as
@@
The set $A^{*}$ is a corner set as in Section 3 of [39] with cardinality $m={l+d\choose l}$ and is unisolvent. Denote the minimal singular value of $V_{A^{*}}$ by $c^{*}$. Then for all $u\in\mathbb{R}^{m}$ with $\|u\|_{2}=1$,
@@
$\sup_{x\in\Omega}|v_{x}^{T}u|\geq c^{*}/\sqrt{m}.$ (42)
Next, for the Vandermonde matrix $V_{A}$ associated with the set $A$, consider its SVD decomposition $V_{A}=U_{L}DU_{R}^{T}$, where both $U_{L}$, $U_{R}$ are unitary matrices and $D$ is a diagonal matrix with positive diagonal entries. Let us denote the left and right singular vectors corresponding to the minimal singular value of $V_{A}$ by $u_{l,m}$ and $u_{r,m}$, respectively. Let $P^{*}$ be the interpolation polynomial such that $P^{*}(A)=u_{r,m}$, then $P^{*}(x)=P^{*}(A)^{T}V_{A}^{-1}v_{x}$. By the lower bound (42),
@@
$\sup_{x\in\Omega}|P^{*}(x)|\leq c_{N}\sup_{x^{\prime}\in A}|P^{*}(x^{\prime})|=c_{N}\|u_{r,m}\|_{\infty}\leq c_{N}.$ (44)
Combining equations (A.2) and (44), we have
@@
A.3 Proof of Lemma 3
@@
In view of Lemma 2, the Vandermonde matrix $V_{A}$ is invertible. Thus following from Lemma 1 all polynomials $P\in\mathscr{P}_{l}(\Omega)$ can be written in the form $P(x)=P(A)^{T}V_{A}^{-1}v_{x}$, $\forall x\in\Omega$, hence
$|P(x)|\leq\|P(A)\|_{\infty}\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1},$
implying the bound $c_{N}\leq\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}.$ On the other hand, since $\Omega$ is a compact set, $\exists x_{0}\in\Omega$ satisfying $\|V_{A}^{-1}v_{x_{0}}\|_{1}=\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}$. Let us take the unique polynomial $P\in\mathscr{P}_{l}(\Omega)$ satisfying $P(A)[i]=(\mathrm{sign}(V_{A}^{-1}v_{x_{0}}))[i]$. Then
@@
This provides the lower bound $c_{N}\geq\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}$ and hence verifies (17).
Next consider a scale parameter $\tau>0$. Let us define the diagonal matrix $\Gamma\in\mathbb{R}^{m\times m}$ with diagonal elements $\Gamma[j,j]=\tau^{|k_{(j)}|}$, where $k_{(j)}$ is the $j$th multi-index in $\mathbb{N}^{d}$ with respect to the lexicographical ordering. Note that $V_{\tau A}=\Gamma V_{A}$ and $v_{\tau x}=\Gamma v_{x}$. Thus the Vandermonde matrix $V_{\tau A}$ is also invertible and $\forall\tau x\in\tau\Omega$, $V_{\tau A}^{-1}v_{\tau x}=V_{A}^{-1}v_{x}$. Therefore, $\tau A$ is a norming set of $\mathscr{P}_{l}(\tau\Omega)$ with the same norming constant $c_{N}$. ∎
A.4 Algorithms for building Layered Norming DAGs
We provide Algorithm 1 used to construct the Layered Norming DAGs discussed in Section 4.2.
1 Input: grid data $\mathcal{X}_{n}$ with $n=\tilde{n}^{d}$ and order $l$. Set $m={l+d\choose l}$ and let $r$ be the largest integer such that $\tilde{n}\geq 2^{r}+1$.
@@ A.4 Algorithms for building Layered Norming DAGs
15         Order the projection of $\cup_{0\leq j^{\prime}\leq j-1}\mathcal{N}_{j^{\prime}}$ onto the $h$th coordinate by the increasing distance to $X_{i}[h]$ as $B_{h}=(x_{s(1),h},x_{s(2),h},\cdots,x_{s(n^{\prime}),h})$.
@@
We briefly discuss the computational complexity of Algorithm 1. We first deal with the case where the dataset $\mathcal{D}_{n}$ is stored in a structured way such that it takes $O(1)$ time to retrieve the data on any grid location. The grid structure ensures a simple, explicit form for the layers $\mathcal{N}_{j}$, without the use of costly numerical methods. Thus the computational complexity of building the layers is $O(n)$. The allocation of parent sets requires sorting with respect to each coordinate. However, due to the grid structure the set $B_{h},\forall 1\leq h\leq d,\forall X_{i}$ has a simple, closed form solution. Therefore, for all $X_{i}$, the allocation of the corresponding parent set, with cardinality $m$, only takes $O(m)$ steps. Overall, the computational complexity of Algorithm 1 is $O(nm)$, with $m={\underline{\alpha}+d\choose\underline{\alpha}}$. If the dataset $\mathcal{D}_{n}$ is not stored in a structured way, one has to start with appropriately sorting it, raising the overall computational complexity to $O(dn\ln n+nm)$.
A.5 Building DAGs satisfying the Norming Condition 2 for general $\mathcal{X}_{n}$
In this section, we discuss how to build DAGs satisfying Conditions 1 and 2 for general design points $\mathcal{X}_{n}$. Without loss of generality, we assume that $\mathcal{X}_{n}\subset[0,1]^{d}$.
The construction of the DAG follows a three-steps procedure. First, we sort all elements in $\mathcal{X}_{n}$ with respect to maximin ordering, defined in a sequential manner. We start by selecting the first element arbitrarily. Once the first $i$ elements are found, the $(i+1)$th element is chosen as the element in $\mathcal{X}_{n}$ that maximizes the minimal distance to the first $i$ elements, with respect to a metric specified by the user. The procedure is repeated until all $n$ elements are sorted. With a slight abuse of notation, we denote the sorted elements as $X_{1},X_{2},\cdots,X_{n}$.
Second, for all $i\geq 2$, let us denote the distance between $X_{i}$ and $\{X_{i^{\prime}},1\leq i^{\prime}<i\}$ as $\mathrm{dist}_{i}$. Let $\mathcal{N}_{0}=\{X_{1}\}$. Then, for all $j\geq 1$, define $\mathcal{N}_{j}$ as
@@ A.5 Building DAGs satisfying the Norming Condition 2 for general $\mathcal{X}_{n}$
for some given constant $\gamma>1$.
Finally, once the layers are specified, for all $j$ large enough such that $\cup_{0\leq j^{\prime}\leq j-1}\mathcal{N}_{j^{\prime}}$ has at least $l^{d}$ elements, for all $X_{i}\in\mathcal{N}_{j}$, we choose the corresponding parent set $\mathrm{pa}(X_{i})$ via the following iterative procedure. Let $c_{S}^{*}>0$ be a given constant and set $c_{S}=c_{S}^{*}$. Furthermore, let use initialize the parent set with the element in $\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$ that is closest to $X_{i}$. Let us then denote the parent set as $\mathrm{pa}(X_{i})=\{w_{1},w_{2},\cdots,w_{m^{\prime}}\}$ after $m^{\prime}\in\{1,...,m-1\}$ iterations. For the next iteration, recall that $k_{(i)},1\leq i\leq m$ denotes the $i$th multi-index with respect to the lexicographical ordering $\prec$” and define the diagonal matrix $\Gamma=\mathrm{diag}\{1,\gamma^{j|k_{(2)}|},\cdots,\gamma^{j|k_{(m)}|}\}\in\mathbb{R}^{m\times m}$. Then, we add the element $X^{\prime}$ to $\mathrm{pa}(X_{i})$ if the matrix
$M=\Gamma\;[v_{w_{1}-X_{i}}\;v_{w_{2}-X_{i}}\cdots v_{w_{m^{\prime}}-X_{i}}\;v_{X^{\prime}-X_{i}}]$ (46)
[…]
1 Input: data set $\mathcal{X}_{n}$, initial singular value $c_{S}^{*}>0$, constant $\gamma>1$ for the layers, order $l$ for polynomials and numerical tolerance $\epsilon_{\mathrm{tol}}$. Set $m={l+d\choose l}$.
@@
3 Set layers $\mathcal{N}_{0}=\{X_{1}\}$ and $\mathcal{N}_{1},\mathcal{N}_{2},\cdots\mathcal{N}_{J}$ according to (45).
@@
10      Sort the elements of $\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$ by increasing distance to $X_{i}$ as $B=(X_{i_{1}},X_{i_{2}},\cdots,X_{i_{n^{\prime}}})$, where $n^{\prime}=|\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}|$.
@@
14           if  $\lambda_{\min}(M)\geq c_{S}$, for matrix $M$ given in (46)

### v4

@@ B.1 Vandermonde matrices and polynomial interpolation
We prove the equivalence statement by contradiction. Suppose the Vandermonde matrix $V_{A}$ is not invertible. Since $V_{A}$ is a square matrix, there exists a nonzero vector $b$, such that $V_{A}b=0$. Assume furthermore, that there exists a polynomial $P(x)$ satisfying ${P}(w_{i})=y_{i},\forall 1\leq i\leq m$; otherwise, the statement holds. Then the polynomial $\tilde{P}(x)=P(x)+\sum_{i=1}^{m}b[i]x^{k_{(i)}}$ also satisfies $\tilde{P}(w_{i})=y_{i},\forall 1\leq i\leq m$, and thus, in view of $b\neq 0$, there exist at least two polynomails satisfying the interpolation condition. This contradicts uniqueness, proving the equivalence.
Next, assume that the Vandermonde matrix $V_{A}$ is invertible. Then, equation (B.3) provides a solution. We prove by contradiction that this solution is unique. Suppose that there are at least two distinct polynomials $P_{1}\neq P_{2}$ satisfying $P_{j}(w_{i})=y_{i}$ for $j=1,2$. Then $P_{1}-P_{2}\neq 0$ and $(P_{1}-P_{2})(w_{i})=0,\forall 1\leq i\leq m$. Writing the difference in the form $(P_{1}-P_{2})(x)=\sum_{i=1}^{m}b[i]x^{k_{(i)}}$ and introducing the notation $b=(b[1],b[2],\cdots,b[m])^{T}\in\mathbb{R}^{m}$, $b\neq 0$, we obtain $V_{A}b=0$, which contradicts with the assumption that $V_{A}$ is invertible. Hence the solution is unique. ∎
B.2 Results regarding Norming sets
In view of Lemma B.1, the unisolvency of the polynomial interpolation problem is equivalent to the invertibility of the corresponding Vandermonde matrix. Building on this, and recalling the notion of norming sets introduced in Section 3.1, we aim to establish qualitative relationships among the properties of the polynomials, the minimal singular value of the Vandermonde matrices, and the spatial dispersion of the set $A$, beginning with the following lemma.
Lemma B.2.
Let $\Omega=[0,1]^{d}$. For $l\in\mathbb{N}$, and a set $A=\{w_{1},w_{2},\cdots,w_{m}\}$ with cardinality $m={l+d\choose l}$, the following two statements are equivalent:
  1. (1)
    The set $A$ is a norming set for $\mathscr{P}_{l}(\Omega)$ with norming constant $c_{N}>0$;
  2. (2)
    The minimal singular value of the Vandermonde matrix $V_{A}$ is bounded from below by $c_{S}>0$.
Moreover, there exists a universal constant $C_{0}\geq 1$, independent of $A$, such that
$C_{0}^{-1}c_{N}\leq c_{S}^{-1}\leq C_{0}c_{N}.$
@@ B.2 Results regarding Norming sets
We first prove (2) $\Rightarrow$ (1). Since the smallest singular value of the Vandermonde matrix $V_{A}$ is positive, $V_{A}$ is invertible. In view of Lemma B.1, the set $A$ is unisolvent with respect to $\mathscr{P}_{l}(\mathbb{R}^{d})$ and, therefore, also unisolvent for $\mathscr{P}_{l}(\Omega)$. Consequently, every polynomial $P\in\mathscr{P}_{l}(\mathbb{R}^{d})$ can be expressed in the form
$P(x)=P(A)^{T}V_{A}^{-1}v_{x}.$ (B.4)
@@
We proceed to prove (1) $\Rightarrow$ (2). First we verify by contradiction that the Vandermonde matrix $V_{A}$ is invertible. Suppose it is not. Then there exists a non-zero vector $b\in\mathbb{R}^{m}$ such that $V_{A}^{T}b=0$. The corresponding polynomial $P(x)=b^{T}v_{x}\in\mathcal{P}_{l}(\Omega)$ satisfies $P(x)=0$, $\forall x\in A$, i.e. $\|P\|_{A}=0$. However, since $b\neq 0$, $\exists x^{\prime}\in\Omega$ such that $P(x^{\prime})\neq 0$, i.e. $\sup_{x}|P(x)|>0$. This contradicts the assumption that $A$ is a norming set.
To establish $c_{S}^{-1}<C_{0}c_{N}$, we introduce a “standard” norming set $A^{*}$ to serve as a “bridge” between $c_{S}$ and $c_{N}$. Define
@@
The set $A^{*}$ is a corner set, as in Section 3 of [40] with cardinality $m={l+d\choose l}$, and is unisolvent. Denote the minimal singular value of $V_{A^{*}}$ by $c^{*}$. Then for all $u\in\mathbb{R}^{m}$ with $\|u\|_{2}=1$,
@@
$\sup_{x\in\Omega}|v_{x}^{T}u|\geq c^{*}/\sqrt{m}.$ (B.5)
Next, for the Vandermonde matrix $V_{A}$ associated with the set $A$, consider its singular value decomposition (SVD) $V_{A}=U_{L}DU_{R}^{T}$, where both $U_{L}$, $U_{R}$ are unitary matrices, and $D$ is a diagonal matrix with positive diagonal entries. Let $u_{l,m}$ and $u_{r,m}$ denote the left and right singular vectors corresponding to the minimal singular value of $V_{A}$, respectively. Let $P^{*}$ be the interpolation polynomial such that $P^{*}(A)=u_{r,m}$. Then $P^{*}(x)=P^{*}(A)^{T}V_{A}^{-1}v_{x}$. By the lower bound in (B.5),
@@
$\sup_{x\in\Omega}|P^{*}(x)|\leq c_{N}\sup_{x^{\prime}\in A}|P^{*}(x^{\prime})|=c_{N}\|u_{r,m}\|_{\infty}\leq c_{N}.$ (B.7)
Combining equations (B.2) and (B.7), we have
@@
Since the constant $c^{*}$ is independent of the set $A$, we finished the proof. ∎
While the equivalence between norming sets and unisolvent sets is well known in the literature, the above lemma characterizes the qualitative relationship between the constants $c_{N}$ and $c_{S}$. This explicit dependence plays a crucial role in computing the small deviation bounds for Vecchia GPs in Section 4. We note, that an important difference between Lemma B.1 and Lemma B.2 lies in the support of the polynomials. In view of the unboundedness of polynomials on the whole domain $\mathbb{R}^{d}$, finite norming sets do not exist when $\Omega=\mathbb{R}^{d}$.
The norming constant $c_{N}$ has an explicit expression and is “scale-invariant” as stated in the following lemma, which is also discussed in [8].
Lemma B.3.
Let $\Omega$ be a compact subset of $\mathbb{R}^{d}$ and let $A\subset\Omega$ with cardinality $m={l+d\choose l}$ be a norming set for $\mathscr{P}_{l}(\Omega)$ with norming constant $c_{N}$. Then we have
$c_{N}=\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}.$ (B.8)
Moreover, for all $\tau>0$, the set $\tau A$ is a norming set for $\mathscr{P}_{l}(\tau\Omega)$ with the same norming constant $c_{N}$.
@@
In view of Lemma B.2, the Vandermonde matrix $V_{A}$ is invertible. Thus following from Lemma B.1, every polynomial $P\in\mathscr{P}_{l}(\Omega)$ can be expressed as $P(x)=P(A)^{T}V_{A}^{-1}v_{x}$, $\forall x\in\Omega$, which implies
$|P(x)|\leq\|P(A)\|_{\infty}\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}.$
Hence, we obtain the bound $c_{N}\leq\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}.$ On the other hand, since $\Omega$ is compact, $\exists x_{0}\in\Omega$ satisfies $\|V_{A}^{-1}v_{x_{0}}\|_{1}=\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}$. Let us take the unique polynomial $P\in\mathscr{P}_{l}(\Omega)$ satisfying $P(A)[i]=(\mathrm{sign}(V_{A}^{-1}v_{x_{0}}))[i]$. Then
@@
This provides the lower bound $c_{N}\geq\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}$ and hence verifies (B.8).
Next, consider a scale parameter $\tau>0$. Define the diagonal matrix $\Gamma\in\mathbb{R}^{m\times m}$ with diagonal entries $\Gamma[j,j]=\tau^{|k_{(j)}|}$, where $k_{(j)}$ is the $j$th multi-index in $\mathbb{N}^{d}$ with respect to lexicographical ordering. Note that $V_{\tau A}=\Gamma V_{A}$ and $v_{\tau x}=\Gamma v_{x}$. Thus the Vandermonde matrix $V_{\tau A}$ is also invertible, and $\forall\tau x\in\tau\Omega$, $V_{\tau A}^{-1}v_{\tau x}=V_{A}^{-1}v_{x}$. Therefore, the scaled set $\tau A$ is a norming set for $\mathscr{P}_{l}(\tau\Omega)$ with the same norming constant $c_{N}$. ∎
While the above lemma provides the explicit formula (B.8) for the norming constant $c_{N}$, computing it can still be challenging, since it requires inverting the corresponding Vandermonde matrix and taking a supremum over the entire compact set $\Omega$. In the unidimensional case, $\Omega=[0,1]$, we can provide an alternative formulation. Let $A=\{w_{1},w_{2},\cdots,w_{l+1}\}\subset[0,1]$ consist of $l+1$ distinct points. Then $c_{N}$ can be expressed using Lagrange polynomials as
$c_{N}=\sup_{x\in[0,1]}\sum_{j=1}^{l+1}\bigg|\frac{\prod_{i\neq j}(x-w_{i})}{\prod_{i\neq j}(w_{j}-w_{i})}\bigg|.$ (B.9)
In higher dimensions, however, without additional technical assumptions on the set $A$, equation (B.8) can not be further simplified.
[…]
First, note that as the polynomial order $l$, and consequently the cardinality of the corresponding norming sets, increases, the norming constants also increase. This is intuitive: as the dimension of the polynomial space grows with $l$, it becomes more difficult to characterize all polynomials using a finite set of functionals. Second, one can observe that for randomly chosen points $w_{j}$, the norming constant is substantially larger than for the equidistant design.
Next, we consider the case $d=2$, see Figure B.2. Similarly to the one-dimensional case,the norming constants increase with the order of the polynomials. Furthermore, the norming constants are smaller for the so called corner sets than for i.i.d. samples generated from a uniform distribution. A corner set is a carefully designed unisolvent set that captures the local polynomial structure, see Section 3.3 for its definition and further discussion. Moreover, for $d\geq 2$, unlike in the one-dimensional case, not every set with cardinality $l+d\choose l$} is a norming set for $\mathscr{P}_{l}(\Omega)$. Examples of such non-norming sets are shown in the bottom row of Figure B.2.
Figure B.1: Norming sets for a one-dimensional space for polynomials of order $1,2,3$ and $4$, respectively. Top row: equally spaced points on $[0,1]$; bottom row: i.i.d. samples from the uniform distribution on $[0,1]$.
Figure B.2: Norming sets in $0,1]^{2}$. The cardinality of the sets in the $l$th column ($1\leq l\leq 4$) is $l+2\choose l$}, which equals the dimension of the polynomial space $\mathscr{P}_{l}(\Omega)$. Top row: corner sets; middle row: i.i.d. samples from the uniform distribution; bottom row: non-norming sets for $\mathscr{P}_{l}(\Omega)$.
C Algorithms for building layered norming DAGs
C.1 Algorithms for building layered norming DAGs
We provide Algorithm 1 used to construct the layered norming DAGs discussed in Section $3.$
1Input: grid data $\mathcal{X}_{n}$ with $n=\tilde{n}^{d}$ and order $l$. Set $m={l+d\choose l}$ and let $r$ be the largest integer such that $\tilde{n}\geq 2^{r}+1$.
@@ C.1 Algorithms for building layered norming DAGs
15         Order the projection of $\cup_{0\leq j^{\prime}\leq j-1}\mathcal{N}_{j^{\prime}}$ onto the $h$th coordinate by increasing distance to $X_{i}[h]$ as $B_{h}=(x_{s(1),h},x_{s(2),h},\cdots,x_{s(n^{\prime}),h})$.
@@
We briefly discuss the computational complexity of Algorithm 1. We first consider the case where the dataset $\mathcal{D}_{n}$ is stored in a structured format, so that retrieving data at any grid point takes $O(1)$ time. The grid structure ensures a simple, explicit form for the layers $\mathcal{N}_{j}$, without requiring costly numerical methods. Thus the computational complexity of building the layers is $O(n)$. Allocating parent sets requires sorting with respect to each coordinate. However, due to the grid structure, the set $B_{h},\forall 1\leq h\leq d$ and all $X_{i}$ has a simple, closed-form solution. Therefore, for each $X_{i}$, assigning its parent set of cardinality $m$ requires only $O(m)$ steps. Overall, the computational complexity of Algorithm 1 is $O(nm)$, with $m={\underline{\alpha}+d\choose\underline{\alpha}}$. If the dataset $\mathcal{D}_{n}$ is not stored in a structured format, one has to start with appropriately sorting it, increasing the overall computational complexity to $O(dn\ln n+nm)$.
C.2 Building DAGs satisfying the Norming Condition 2 for general $\mathcal{X}_{n}$
In this section, we discuss how to build DAGs satisfying Conditions 1 and 2 for general design points $\mathcal{X}_{n}$. Without loss of generality, we assume $\mathcal{X}_{n}\subset[0,1]^{d}$.
The DAG construction follows a three-steps procedure. First, we sort all elements in $\mathcal{X}_{n}$ according to maximin ordering, defined sequentially. The first element is selected arbitrarily. Once the first $i$ elements are chosen, the $(i+1)$th element is selected as the point in $\mathcal{X}_{n}$ that maximizes the minimal distance to the first $i$ elements, with respect to a metric specified by the user. The procedure is repeated until all $n$ elements are sorted. With a slight abuse of notation, we denote the sorted elements as $X_{1},X_{2},\cdots,X_{n}$.
Second, for all $i\geq 2$, let $\mathrm{dist}_{i}$ denote the distance between $X_{i}$ and the set $\{X_{i^{\prime}},1\leq i^{\prime}<i\}$ as . Let $\mathcal{N}_{0}=\{X_{1}\}$. Then, for all $j\geq 1$, define the layers $\mathcal{N}_{j}$ as
@@ C.2 Building DAGs satisfying the Norming Condition 2 for general $\mathcal{X}_{n}$
for a given constant $\gamma>1$.
Finally, once the layers are specified, for all $j$ large enough such that $\cup_{0\leq j^{\prime}\leq j-1}\mathcal{N}_{j^{\prime}}$ contains at least $l^{d}$ elements, for each $X_{i}\in\mathcal{N}_{j}$, we choose the corresponding parent set $\mathrm{pa}(X_{i})$ via the following iterative procedure. Let $c_{S}^{*}>0$ be a given constant and set $c_{S}=c_{S}^{*}$. Initialize $\mathrm{pa}(X_{i})$ with the element in $\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$ that is closest to $X_{i}$. Suppose that after $m^{\prime}\in\{1,...,m-1\}$ iterations, the parent set is $\mathrm{pa}(X_{i})=\{w_{1},w_{2},\cdots,w_{m^{\prime}}\}$. Recall that $k_{(i)},1\leq i\leq m$, denotes the $i$th multi-index in lexicographical ordering $\prec$” and define the diagonal matrix $\Gamma=\mathrm{diag}\{1,\gamma^{j|k_{(2)}|},\cdots,\gamma^{j|k_{(m)}|}\}\in\mathbb{R}^{m\times m}$. Then, we add a new element $X^{\prime}$ to $\mathrm{pa}(X_{i})$ if the matrix
$M=\Gamma\;[v_{w_{1}-X_{i}}\;v_{w_{2}-X_{i}}\cdots v_{w_{m^{\prime}}-X_{i}}\;v_{X^{\prime}-X_{i}}]$ (C.2)
satisfies $\sigma_{\min}(M)\geq c_{S}$, where $\sigma_{\min}(M)$ denotes the minimum singular value of $M$, and $v_{x}$ is defined in (B.2). If all $m$ elements are selected this manner, the process terminates. Otherwise, we set $c_{S}=c_{S}/2$ and repeat the procedure with this updated value. The process terminates either when a parent set of size $m$ is constructed, or when $c_{S}$ drops below a prescribed, numerical tolerance level. In the latter case, we simply fill up the parent set with the closest data points to $X_{i}$ until the prescribed cardinality $m$ is reached. We formalize this procedure in Algorithm 2.
1Input: data set $\mathcal{X}_{n}$, initial singular value $c_{S}^{*}>0$, constant $\gamma>1$ for the layers, order $l$ for polynomials and numerical tolerance $\epsilon_{\mathrm{tol}}$. Set $m={l+d\choose l}$.
@@
3 Set layers $\mathcal{N}_{0}=\{X_{1}\}$ and $\mathcal{N}_{1},\mathcal{N}_{2},\cdots\mathcal{N}_{J}$ according to (C.1).
@@
10      Sort the elements of $\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$ in order of increasing distance to $X_{i}$, denoted as $B=(X_{i_{1}},X_{i_{2}},\cdots,X_{i_{n^{\prime}}})$, where $n^{\prime}=|\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}|$.
@@
14           if $\lambda_{\min}(M)\geq c_{S}$, for matrix $M$ given in (C.2)